7. Find the 31st term of an A.P. whose 11th term is 38 and the 16th term is 73.

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Given that, 11th term, a11 = 38 and 16th term, a16 = 73 We know that, an = a+(n−1)d a11 = a+(11−1)d 38 = a+10d ………………………………. (i) In the same way, a16 = a +(16−1)d 73 = a+15d ………………………………………… (ii) On subtracting equation (i) from (ii), we get 35 = 5d d = 7 From equation (i), we can write, 38 = a+10×(7) 38 − 70 = a a = −32 a31 = a +(31−1) d = − 32 + 30 (7) = − 32 + 210 = 178 Hence, 31st term is 178.